OpenStudy (ohmath123):

4^3x-3= 10^7x-10 How can I solve this using logarithms?

jhonyy9 (jhonyy9):

for eaxmple make ln on both sides ln (4^(3x-3) ) = ln (10^(7x-10)) so than what property of logarithms you know ? where go these exponents of 4 and on the other side of 10 ?

OpenStudy (ohmath123):

I would have thought i would put log 4^(3x-3)= log 10^(7x-10) then the log 10 would cancel? @jhonyy9

jhonyy9 (jhonyy9):

no you need to know that log a^x = xlog a

jhonyy9 (jhonyy9):

@Directrix do oyu have other way idea to solve it please ?

jhonyy9 (jhonyy9):

do you - sorry

Directrix (directrix):

I am not sure what the problem states. No grouping symbols.

OpenStudy (ohmath123):

the problem is telling me to solve for x but i have to use logarithms @Directrix @jhonyy9

jhonyy9 (jhonyy9):

you need to solve it for x - but the unknowed x there are in the exponents and this is again mistaque that the base numbers not are the same - so there are 4 and 10

Directrix (directrix):